OpenStudy (anonymous):

( x^3 + x^2 -6x -3 ) / ( x + K ) yields a remainder -3. find the value of k that satisfies the condition p.s. i tried doing synthetic and my remainder is -k^3 + k^2 + 6k -3. so i'm not so sure that helped..

OpenStudy (amistre64):

that does help

OpenStudy (amistre64):

you need to solve that equation for -3 now

Parth (parthkohli):

Yes, yes it does! Equation!

OpenStudy (amistre64):

-k^3 + k^2 + 6k -3 = -3 -k^3 + k^2 + 6k = 0 -k (k^2 - k - 6) = 0

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

then k=0 and , 3 & -2 ?

OpenStudy (amistre64):

those are options to test out, yes. If they all fit the bill; then use them all

OpenStudy (amistre64):

the remainder thrm says, that if you input a value into the function, you get the remainder

OpenStudy (anonymous):

okay, thank you so much ! :)